OpenStudy (anonymous):

MEDAL ASAP what are the zeroes of the function y=(x-2)(x-3)(x+3)

OpenStudy (amistre64):

the values that make y=0 of course ....

OpenStudy (amistre64):

you may want to review the zero product rule if this isnt making sense

OpenStudy (anonymous):

I got -2,3,-3 but I am not sure if that is correct.

OpenStudy (kirbykirby):

Almost you need: (x-2)=0 (x-3)=0 (x+3)=0 (check the first equation)

OpenStudy (anonymous):

ok so now i have 2,3,-3 haha

OpenStudy (kirbykirby):

yep :)